Luke Company has three divisions: Peak, View, and Grand. The company has a hurdle rate of 6.76 percent. Selected operating data for the three divisions follow:  

Peak View Grand
Sales revenue $ 339,000 $ 236,000 $ 299,000
Cost of goods sold 210,000 120,000 193,000
Miscellaneous operating expenses 42,000 39,000 33,000
Average invested assets 1,320,000 930,000 1,025,000


Required:
1. Compute the return on investment for each division. (Enter your ROI answers as a percentage rounded to two decimal places, (i.e., 0.1234 should be entered as 12.34%.))

       

2. Compute the residual income for each division. (Loss amounts should be indicated by a minus sign. Round your answers to nearest whole dollar.)
Answer #1

Answer:

Return on Investment = (Sales Revenue - COGS - Misc Operating Expenses)/Average Invested Assets

Peak = (339000-(210000+42000)) / 1320000 = 6.59%

View = (236000-(120000+39000)) / 930000 = 8.28%

Grand = (299000-(193000+33000)) / 1025000 = 7.12%

Residual Income = Operating Income - (Average invested assets X Hurdle Rate)

Peak = 87000 - (1320000 X 6.76%) = -2,232

View = 77000 - (930000 X 6.76%) = 14,132

Grand = 73000 - (1025000 X 6.76%) = 3,710